Shelby has some nickels and quarters worth $2.30. Andrew has twice as many nickels as Shelby, but 3 less quarters that Shelby. Andrew's coins are worth $1.85. How many nickels and quarters does Shelby have? Any help is appreciated!
Answer by ikleyn(52776)

.

Let N and Q are the numbers of nickels and quarters that Shelby has (the values under the question).


Then Andrew has 2N nickels and (Q-3) quarters.


You have these 2 equations for 2 unknowns


    10N     + 25Q      = 230  cents   (1)    (Shelby's money)

    10*(2N) + 25*(Q-3) = 185 cents    (2)    (Andrew's money)


Simplify and reduce to the standard form


    10N + 25Q = 230     (3)

    20N + 25Q = 260     (4)


From equation (4), subtract equation (3)  (both sides)


     10N      = 30


Hence,  N = 3.


Then from equation (1),  


    10*3 + 25Q = 230,

    25Q = 230 - 30 = 200

      Q            = 200/25 = 8.


ANSWER.  Shelby has  3 nickels and 8 quarters.

You may check that my answer satisfies all conditions of the problem.

Do it on your own (!)